# Virne: A Comprehensive Benchmark for Deep RL-based Network Resource Allocation in NFV

### v1

@@ 3.3.2 Unified Pipeline for Efficient Implementations
Considering this widely-used MDP model, we generally unify existing NFV-RA RL-based algorithms into three key components: MDP modeling, policy architecture, and training methods. These methods model the NFV-RA solution construction process as above-introduced MDPs with a specific reward design and feature engineering. Subsequently, they use various neural networks to build policy architectures, such as Convolutional Neural Networks (CNN) nn-nips-2012-cnn and Graph Convolutional Network (GCN) gnn-iclr-2017-gcn . These policies are trained with a selected RL method, such as Asynchronous Advantage Actor-critic (A3C) rl-icml-2016-a3c and Proximal Policy Optimization (PPO) drl-arxiv-2017-ppo .
@@ 3.3.3 Implemented RL-based NFV-RA Algorithms
To provide clarity in both implementation and evaluation, we categorize these works under common names that reflect RL algorithms and their neural policy architectures. To further streamline the training process and avoid unnecessary inconsistencies, we adopt PPO as the default RL algorithm in our experiments, due to its recognized training efficiency and strong empirical performance drl-tnnls-2022-drl-survey . For example, the PPO-GCN method uses graph convolutional networks as feature encoders while employing Proximal Policy Optimization for efficient training. In the subsequent experiments, we will first investigate the impact of distinct implementation techniques, such as reward function and feature engineering, and identify the most efficient configurations for general implementation. Then, all of them will be evaluated for their effectiveness in the cloud and other popular scenarios.
@@ 3.5 Evaluation Criteria
Performance Metrics. Virne includes the critical performance metrics of NFV-RA algorithms vne-cst-2013-survey-vne ; nfv-comst-2024-ai-vne-survey , including request acceptance rate (RAC), long-term revenue-to-cost (LRC), long-term average revenue (LAR) and average solving time (AST). See Appendix C.4 for their definitions.
Practicality Perspectives. To comprehensively assess the practicality of NFV-RA algorithms, we examine multiple perspectives that extend beyond mere effectiveness. In particular, we consider three aspects of NFV-RA algorithms: (a) Solvability denotes the ability to find feasible solutions; (b) Generalization indicates reliable performance across various network conditions; (c) Scalability measures how effectively it accommodates increases in network size and complexity. Gaining these insights with our pre-provided evaluation interfaces, Virne helps users understand the practical viability of an algorithm and guides further development and application.
@@ 4.1 Experimental Setup
Simulation Settings. We adopt the widely-used topologies as physical networks: synthetic WX100 vne-dataset-jsac-1988-waxman , and real-world GEANT and BRAIN vne-network-2010-dataset-sdnlib , which cover various network scales and densities. See Appendix D.1.2 for their detailed descriptions. Computing resources of physical nodes and bandwidth resources of physical links are uniformly distributed within the same range of units, i.e., $\mathcal{X}_{C(n_{p})}\sim\mathcal{U}(50,100)$ and $\mathcal{X}_{B(l_{p})}\sim\mathcal{U}(50,100)$. In default settings, for each simulation run, we create 1000 VN with varying sizes following a uniform distribution, $\mathcal{X}_{|\mathcal{G}_{v}|}\sim\mathcal{U}(2,10)$. The computing resource demands of the nodes and the bandwidth requirements of the links within each VN are uniformly distributed, i.e., $\mathcal{X}_{C(n_{v})}\sim\mathcal{U}(0,20)$ and $\mathcal{X}_{B(l_{v})}\sim\mathcal{U}(0,50)$, respectively. The virtual nodes in each VN are randomly interconnected with a probability of 50%. The lifetime of each VN follows an exponential distribution with an average of 500 time units. The arrival of these VNs follows a Poisson process with an average rate $\eta$, where $\eta$ denotes the average arrived VN count per unit of time. Due to the varying physical resource supply in these topologies caused by distinct scale and density, we use different $\eta$, i.e., (0.16, 0.016, 0.004) for (WX100, GEANT, BRAIN), to accommodate the reasonable request demand. In the subsequent experiments, we manipulate the distribution settings of VNs and change the PN topologies to simulate various network systems.
@@ 4.2.1 Exploration on Implementation Techniques
Variations in core implementation techniques, such as reward function designs, feature engineering, action masking strategies, lead to substantial differences in the performance of RL-based NFV-RA algorithms. We study their impacts by systematically evaluating three representative policy architectures (PPO-MLP, PPO-ATT, PPO-DualGAT) in WX100, summarized in Table 2.
@@
Impact on Reward Function Design. The design and magnitude of intermediate rewards strongly affect training and final performance. Across all solvers, a moderate fixed intermediate reward (fixed, 0.1) consistently yields the best or near-best results. In contrast, very small or no intermediate rewards result in clear performance drops, highlighting the importance of sufficient reward signals for effective exploration. Interestingly, the adaptive intermediate reward, which intuitively normalizes the total reward for different VN sizes, does not achieve optimal performance in practice. This suggests that, 1 while normalization is conceptually appealing, a fixed and moderate reward signal is more effective for guiding policy learning in the NFV-RA context.
@@
Impact on Feature Engineering Combination. We compare the use of Status (S), Topological (T), and their combination in the feature constructor. The results show that combining both Status and Topological features (✓, ✓) generally leads to the best or second-best outcomes for acceptance rate and revenue, regardless of the solver architecture. For example, PPO-ATT+ achieves the best acceptance rate (0.712) and LRC (0.661) with both features, and a marked drop when using only one. This demonstrates that 2 topological metrics serve as a valuable augmentation, capturing global network context and node importance, even for graph neural networks such as GATs.
@@
Impact on Action Mask Strategy Applying action masking (✓), which prevents the agent from selecting infeasible actions (e.g., resource availability satisfaction), consistently improves performance across all methods. For instance, in PPO-MLP+ and PPO-DualGAT+, enabling action masking increases acceptance rate by up to 0.053 compared to otherwise identical configurations without masking. This demonstrates that 3 explicit constraint enforcement is vital for robust RL-based NFV-RA solutions, since the constraints of NFV-RA are intricate and hard.
@@ 4.2.2 Effectiveness in Online Environments
To evaluate the effectiveness of these NFV-RA algorithms, we conduct online simulations across three distinct network topologies: WX100, GEANT, and BRAIN, each with different traffic rates ($\eta$) as specified. Table 3 reports experimental results of the implemented RL-based and traditional algorithms on key metrics. Particularly, RL agents equipped with dual graph neural network architectures, such as PPO-DualGAT and PPO-DualGCN, consistently dominate across diverse topologies (WX100, GEANT, BRAIN). PPO-DualGAT, for instance, achieves a remarkable 78.10% RAC and the highest LAR (14938.60) in WX100, while PPO-DualGCN excels in the GEANT network with a 60.40% RAC and 0.75 LRC. This suggests that 4 the ability of these dual-GNN models to concurrently process and relate features from both the virtual and physical network graphs provides a significant edge in finding high-quality embeddings. Traditional heuristics, though extremely fast (e.g., GRC-Rank with AST $\sim$0.01-0.04s), generally lag in solution quality (e.g., GRC-Rank LRC 0.56 in WX100 vs. PPO-DualGAT’s 0.74). Meta-heuristics or MCTS can achieve competitive RACs (74.30% for MCTS in WX100), yet their substantial computational overhead (AST >3s for MCTS) renders them less viable for agile, real-time NFV deployments compared to the sub-second inference of RL agents.
4.3 Evaluation from Practicality Perspectives
Generalization on Network Conditions
Network conditions are inherently complex and subject to continuous changes, such as fluctuations in request frequencies and varying resource demands. As such, evaluating the generalization of these trained NFV-RA policies is critical to ensure they can adapt effectively to different, evolving network environments. To address this, we conduct the experiments in various network conditions to study the generalization to conditions of pretrained models, including (a) Evaluation on Varying Traffic Rates and (b) Evaluation on Fluctuating Demand Distribution. See Appendix D.2 for detailed experimental setup and result analysis.
Solvability via Offline Evaluation
Traditional online testing of NFV-RA algorithms, while reflecting real-world scenarios, complicates direct comparisons of algorithmic solvability. The continuous evolution of the PN state, coupled with dynamic VN arrivals and lifetimes, makes it difficult to isolate performance. Specifically, it is challenging to determine if solution failures arise from an algorithm’s inherent limitations or from transient unsolvable network states, thereby hindering fair comparisons on solvability. To address this, Virne provides a controlled offline evaluation environment. This features a series of static instances, each comprising a PN and a VN, evaluating algorithms’ abilities to find high-quality, feasible solutions under reproducible conditions. Through such evaluation, we assess the fundamental solvability of each method, not only across an entire dataset but also for distinct VN scales. For detailed experimental setup and results, please see Appendix D.3.
Scalability on Network Sizes
Network systems are growing in size as the physical infrastructure extends and service requests increase. Thus, assessing the scalability of NFV-RA policies is essential to ensure they can maintain their effectiveness as the network size and complexity increase. To assess this, Virne supports scalability evaluations through two primary perspectives: (a) Performance on Large-scaled Networks and (b) analysis of Solving Time Scale as problem size grows. Detailed experimental setups and results for these evaluations are placed in Appendix D.4.
4.4 Validation on Emerging Networks
4.5 Discussion on Future Direction
While RL-based methods for NFV-RA show promise, there are still practical challenges through our empirical observation. To advance deep RL for NFV-RA, we identify several future directions that also pose significant on existing ML methods. They are included but not limited to (a) develop sophisticated representation learning for cross-graph statuses and attributed constraints, (b) achieving generalizable policies adaptable to varying network scales and dynamic conditions, (c) creating robust learning frameworks to enable learning in conflicting operational constraints, and (d) engineering highly scalable algorithms for extremely large-scale physical infrastructures. A more detailed exploration of these prospective is provided in the supplementary material.
@@ A.1.1 Constraint Conditions
$\begin{split}\sum_{n_{p}^{i}\in\Omega(n_{p}^{k})}y^{m,w}_{i,k}-\sum_{n_{p}^{j}\in\Omega(n_{p}^{k})}y^{m,w}_{k,j}=x^{m}_{k}-x^{w}_{k},\forall l^{v}_{m,w}\in\mathcal{L}_{v},n_{v}^{k}\in\mathcal{N}_{p},\end{split}$ (6)
@@
Here, $\Omega(n_{p}^{k})$ denotes the neighbors of $n_{p}^{k}$. Constraint (3) ensures that every virtual node is mapped to one and only one physical node. Conversely, constraint (4 )limits each physical node to hosting at most one virtual node, thus enforcing a unique mapping relationship. Constraint (5) verifies that virtual nodes are allocated to physical nodes with adequate resources. Following the principle of flow conservation, constraint (6) guarantees that each virtual link $(n_{v}^{m},n_{v}^{w})$ is routed along a physical path from $n_{p}^{i}$ (the physical node where $n_{v}^{m}$ is placed) to $n_{p}^{j}$ (the physical node where $n_{v}^{w}$ is placed). Constraint (7) eliminates the possibility of routing loops, thereby ensuring that virtual links are routed acyclically. Lastly, constraint (8) ensures that the bandwidth usage on each physical link remains within its available capacity. Overall, constraints (3,4,5) enforce the one-to-one placement and computing resource availability required in the node mapping process. And constraints (6,7,8) ensure the path connectivity and bandwidth resource availability required in the link mapping process.

### v2

@@ 3.3.2 Unified Pipeline for Efficient Implementations
Considering this widely-used MDP model, we generally unify existing NFV-RA RL-based algorithms into three key components: MDP modeling, policy architecture, and training methods. These methods model the NFV-RA solution construction process as above-introduced MDPs with a specific reward design and feature engineering. Subsequently, they use various neural networks to build policy architectures, such as Convolutional Neural Networks (CNN) (Krizhevsky et al., 2012) and Graph Convolutional Network (GCN) (Kipf and Welling, 2017). These policies are trained with a selected RL method, such as Asynchronous Advantage Actor-critic (A3C) (Mnih et al., 2016) and Proximal Policy Optimization (PPO) (Schulman et al., 2017).
@@ 3.3.3 Implemented RL-based NFV-RA Algorithms
To provide clarity in both implementation and evaluation, we categorize these works under common names that reflect RL algorithms and their neural policy architectures. For example, the PPO-GCN method uses graph convolutional networks as feature encoders while employing Proximal Policy Optimization for efficient training. In the subsequent experiments, we will first investigate the impact of distinct implementation techniques, such as reward function and feature engineering, and identify the most efficient configurations for general implementation. Then, we evaluate their effectiveness in the cloud and other popular scenarios.
@@ 3.5 Evaluation Criteria
Performance Metrics. Virne includes the critical performance metrics of NFV-RA algorithms (Fischer et al., 2013; Wu et al., 2024a), including request acceptance rate (RAC), long-term revenue-to-cost (LRC), long-term average revenue (LAR) and average solving time (AST). See Appendix C.5 for their definitions.
Practicality Perspectives. To comprehensively assess the practicality of NFV-RA algorithms, we develop multiple emerging evaluation protocols that extend beyond mere online effectiveness. In particular, we consider three aspects of NFV-RA algorithms: (a) Solvability denotes the ability to find feasible solutions; (b) Generalization indicates reliable performance across various network conditions; (c) Scalability measures how effectively it accommodates increases in network size and complexity. Gaining these insights with our pre-provided evaluation interfaces, Virne helps users understand the practical viability of an algorithm and guides further development and application.
@@ 4.1 Experimental Setup
Simulation Settings. We adopt the widely-used topologies as physical networks: synthetic WX100 (Waxman, 1988), and real-world GEANT and BRAIN (Orlowski et al., 2010), which cover various network scales and densities. See Appendix D.1.2 for their detailed descriptions. Computing resources of physical nodes and bandwidth resources of physical links are uniformly distributed within the same range of units, i.e., $\mathcal{X}_{C(n_{p})}\sim\mathcal{U}(50,100)$ and $\mathcal{X}_{B(l_{p})}\sim\mathcal{U}(50,100)$. In default settings, for each simulation run, we create 1000 VN with varying sizes following a uniform distribution, $\mathcal{X}_{|\mathcal{G}_{v}|}\sim\mathcal{U}(2,10)$. The computing resource demands of the nodes and the bandwidth requirements of the links within each VN are uniformly distributed, i.e., $\mathcal{X}_{C(n_{v})}\sim\mathcal{U}(0,20)$ and $\mathcal{X}_{B(l_{v})}\sim\mathcal{U}(0,50)$, respectively. The virtual nodes in each VN are randomly interconnected with a probability of 50%. The lifetime of each VN follows an exponential distribution with an average of 500 time units. The arrival of these VNs follows a Poisson process with an average rate $\eta$, where $\eta$ denotes the average arrived VN count per unit of time. Due to the varying physical resource supply in these topologies caused by distinct scale and density, we use different $\eta$, i.e., (0.16, 0.016, 0.004) for (WX100, GEANT, BRAIN), to accommodate the reasonable request demand. In the subsequent experiments, we manipulate the distribution settings of VNs and change the PN topologies to simulate various network systems.
@@ 4.2.1 Exploration on Implementation Techniques
Due to its training efficiency and strong empirical performance (Wang et al., 2022), we adopt PPO as the default RL algorithm in subsequent experiments. This choice is further supported by the study of Impact on RL Training Methods as follows. Furthermore, there are several core implementation choices, such as reward design, feature engineering, and action masking, that substantially affect the performance of RL-based NFV-RA algorithms. To quantify these effects, we systematically evaluate three representative policy architectures (PPO-MLP, PPO-ATT, PPO-DualGAT) on WX100, with results summarized in Table 2. Next, we elaborate on the analysis of each study.
Figure 4: Training curves of PPO-DualGAT and its variations. Each point represents the R2C ratio over all VN Requests within a single training simulation round.
Impact on RL Training Methods. We evaluate the impact of different RL training algorithms by comparing the performance of PG, A3C, and PPO. All three methods are paired with the same PPO-DualGAT policy architecture. As illustrated by the learning curves in Figure 4, PPO demonstrates significantly superior performance. It converges rapidly to the highest average return, stabilizing at a high level of solution quality early in the training process. In contrast, A3C shows slower convergence and reaches a lower performance plateau. PG struggles the most, exhibiting the slowest learning rate and the lowest final performance among the three. These findings highlight PPO’s sample efficiency and robust performance, justifying its selection as the default RL method for our subsequent experiments.
@@
Impact on Reward Function Design. The design and magnitude of intermediate rewards strongly affect training and final performance. Across all solvers, a moderate fixed intermediate reward (fixed, 0.1) consistently yields the best or near-best results. In contrast, very small or no intermediate rewards result in clear performance drops, highlighting the importance of sufficient reward signals for effective exploration. Interestingly, the adaptive intermediate reward, which intuitively normalizes the total reward for different VN sizes, does not achieve optimal performance in practice. As illustrated in Figure 4, we observe that a well-tuned fixed reward of 0.1 has higher convergence than the adaptive reward, which validates our initial performance observations. The adaptive reward is also outperformed by a smaller fixed reward of 0.01, which further demonstrates that it may introduce a noisy signal for the optimization process, leading to sub-optimality. This suggests that, 1 while normalization is conceptually appealing, a fixed and moderate reward signal is more effective for guiding policy learning in the NFV-RA context.
@@
Impact on Feature Engineering Combination. We compare the use of Status (S), Topological (T), and their combination in the feature constructor. The results show that combining both Status and Topological features (✓, ✓) generally leads to the best or second-best outcomes for acceptance rate and revenue, regardless of the solver architecture. For example, PPO-ATT+ achieves the best acceptance rate (0.712) and LRC (0.661) with both features, and a marked drop when using only one. As illustrated in Figure 4, leveraging these advanced features also contributes to faster and higher convergence. This demonstrates that 2 topological metrics serve as a valuable augmentation, capturing global network context and node importance, even for graph neural networks.
@@
Impact on Action Mask Strategy. Applying action masking (✓), which prevents the agent from selecting infeasible actions (e.g., resource availability satisfaction), consistently improves performance across all methods. For instance, in PPO-MLP+ and PPO-DualGAT+, enabling action masking increases acceptance rate by up to 0.053 compared to otherwise identical configurations without masking. This demonstrates that 3 explicit constraint enforcement is vital for robust RL-based NFV-RA solutions, since the constraints of NFV-RA are intricate and hard.
@@ 4.2.2 Effectiveness in Online Environments
To evaluate the effectiveness of these NFV-RA algorithms, we conduct online simulations across three distinct network topologies: WX100, GEANT, and BRAIN, each with different traffic rates ($\eta$) as specified. Table 3 reports experimental results of the implemented RL-based and traditional algorithms on key metrics. While traditional heuristics are exceptionally fast, they fall short in solution quality. Similarly, meta-heuristics, despite achieving competitive acceptance rates, are hindered by significant computational overhead, making them impractical for real-time environments. This highlights that 4 RL-based methods often offer a favorable balance between solution quality and efficiency, making them a compelling choice for NFV-RA. Advanced RL agents, specifically those with dual graph neural network architectures like PPO-DualGAT and PPO-DualGCN, consistently deliver superior performance. This suggests that 5 dual-GNN models’ ability to concurrently process and relate features from both the virtual and physical network graphs provides a significant edge in finding high-quality embeddings. Furthermore, it is notable that 6 the topological characteristics of PN, such as scale and density, impact the relative performance of NFV-RA algorithms, particularly GNN architectures. For instance, on the smaller-scale GEANT and BRAIN, GCN-based methods achieve more competitive and even better performance. This reveals that GATs with an adaptive aggregation mechanism may be more advantageous in denser topologies for prioritizing critical connections, but provide less improvement in sparse topologies.
4.3 Evaluation from Practicality Perspectives
Solvability via Offline Evaluation. Traditional online testing of NFV-RA algorithms, while reflecting real-world scenarios, complicates direct comparisons of algorithmic solvability. The continuous evolution of the PN state, coupled with dynamic VN arrivals and lifetimes, makes it difficult to isolate performance. Specifically, it is challenging to determine if solution failures arise from an algorithm’s inherent limitations or from transient unsolvable network states, thereby hindering fair comparisons on solvability. To address this, Virne provides a controlled offline evaluation environment. This features a series of static instances, each comprising a PN and a VN, evaluating algorithms’ abilities to find high-quality, feasible solutions under reproducible conditions. Through such evaluation, we assess the fundamental solvability of each method, not only across an entire dataset but also for distinct VN scales. This evaluation protocol offers a granular view of algorithmic solvability as VN complexity grows.
Concretely, we evaluate the solution quality, i.e., R2C, achieved by different algorithms for VNs of varying sizes. We conduct experiments on the representative PN, WX100. The results are shown in Figure 5, which presents average solution quality for VN sizes ranging from 2 to 10. There is a generally decreasing trend as the size of the VN increases due to increased complexity. We observe that 7 RL methods with advanced graph representations exhibit superior solvability across most VN sizes. For instance, PPO-DualGAT consistently achieves the highest R2C ratios, particularly for larger VNs. Conversely, simpler RL policies like PPO-MLP perform well on small VNs but degrade noticeably on larger ones due to limited representation ability. This highlights that 8 more powerful foundation policies excel in the vast search space of large VNs, dynamically prioritizing the most critical nodes and links for embedding. Additionally, it is obvious that 9 traditional heuristics and meta-heuristics show mixed results and generally underperform compared to the top RL methods, especially for larger VNs. Particularly, NEA-Rank initially achieves excellent performance for 2-node VNs, but as the VN size increases, it declines and then falls behind the top RL methods.
Generalization on Network Conditions. Network conditions are inherently complex and subject to continuous changes, such as fluctuations in request frequencies and varying resource demands. As such, evaluating the generalization of these trained NFV-RA policies is critical to ensure they can adapt effectively to different, evolving network environments. To address this, we conduct the experiments in various network conditions to study the generalization to conditions of pretrained models, including (a) Evaluation on Varying Traffic Rates and (b) Evaluation on Fluctuating Demand Distribution. See Appendix D.2 for detailed experimental setup and result analysis.
Scalability on Network Sizes. Network systems are growing in size as the physical infrastructure extends and service requests increase. Thus, assessing the scalability of NFV-RA policies is essential to ensure they can maintain their effectiveness as the network size and complexity increase. To assess this, Virne supports scalability evaluations through two primary perspectives: (a) Performance on Large-scaled Networks and (b) analysis of Solving Time Scale as problem size grows. Detailed experimental setups and results for these evaluations are placed in Appendix D.3.
4.4 Validation on Emerging Networks
4.5 Discussion on Future Direction
While RL-based methods for NFV-RA show promise, there are still practical challenges, as observed through our empirical observation. To advance deep RL for NFV-RA, we identify several future directions that also pose significant challenges to existing ML methods. They are included but not limited to (a) developing sophisticated representation learning for cross-graph statuses and attributed constraints, (b) achieving generalizable policies adaptable to varying network scales and dynamic conditions, (c) creating robust learning frameworks to enable learning in conflicting operational constraints, and (d) engineering highly scalable algorithms for extremely large-scale physical infrastructures. We also provide experimental analysis of emerging solutions along these directions, such as safe RL and meta RL-based methods. See Appendix E and Figure 13 for details.
@@ 5 Conclusion
Ethics Statement
This work introduces Virne, a benchmarking framework designed to accelerate scientific research in ML for network optimization. This study does not involve human subjects or sensitive data. All datasets are either synthetically generated or based on publicly available network topologies from SNDlib. As such, we identify no direct negative societal impacts from this research.
Reproducibility Statement
To ensure reproducibility, we have made our benchmark and all necessary resources publicly available. The complete source code for the Virne framework, including all algorithm implementations and simulation environments, is provided in the anonymous codebase link. We provide a detailed description of the experimental setup in Section 4.1 and Appendix D. These details consist of implementation details, network topologies, and data generation. Overall, these resources ensure the full reproducibility of our work and empower the community to build upon our findings.
Acknowledgment
This work was supported in part by the National Key R&D Program of China (Grant No.2023YFF0725001), in part by the National Natural Science Foundation of China (Grant No.92370204), in part by the Guangdong Basic and Applied Basic Research Foundation (Grant No.2023B1515120057), in part by the Key-Area Special Project of Guangdong Provincial Ordinary Universities (2024ZDZX1007).
@@ A.1.1 Constraint Conditions
$\begin{split}\sum_{n_{p}^{i}\in\Omega(n_{p}^{k})}y^{m,w}_{k,j}-\sum_{n_{p}^{j}\in\Omega(n_{p}^{k})}y^{m,w}_{i,k}=x^{m}_{k}-x^{w}_{k},\forall l^{v}_{m,w}\in\mathcal{L}_{v},n_{v}^{k}\in\mathcal{N}_{p},\end{split}$ (5)
@@
Here, $\Omega(n_{p}^{k})$ denotes the neighbors of $n_{p}^{k}$. Constraint (2) ensures that every virtual node is mapped to one and only one physical node. Conversely, constraint (3 )limits each physical node to hosting at most one virtual node, thus enforcing a unique mapping relationship. This one-to-one mapping constraint enforces intra-request node disjointness, a standard requirement in virtual network embedding to ensure survivability and prevent single points of failure (Fischer et al., 2013). Note that this does not preclude inter-request multi-tenancy; virtual nodes from different VNs can share the same physical node, provided resources allow. Constraint (4) verifies that virtual nodes are allocated to physical nodes with adequate resources. Following the principle of flow conservation, constraint (5) guarantees that each virtual link $(n_{v}^{m},n_{v}^{w})$ is routed along a physical path from $n_{p}^{i}$ (the physical node where $n_{v}^{m}$ is placed) to $n_{p}^{j}$ (the physical node where $n_{v}^{w}$ is placed). Constraint (6) eliminates the possibility of routing loops, thereby ensuring that virtual links are routed acyclically. Lastly, constraint (7) ensures that the bandwidth usage on each physical link remains within its available capacity. Overall, constraints (2,3,4) enforce the one-to-one placement and computing resource availability required in the node mapping process. And constraints (5,6,7) ensure the path connectivity and bandwidth resource availability required in the link mapping process.